2. A man who is in the boat at rest, pulls with the help of a rope another boat, which is also at rest, with the force F1 = 800N. The weight of the first boat together with the man is 400kg, while that of the second boat is 320kg. Find: a) The acceleration with which each boat moves. b) The displacements they make during the 3s of the movement. Friction is not taken into account.


1
Expert's answer
2021-02-07T19:10:49-0500

a)


"F_1=-F_2"


"a_1=F_1/m_1=800/400=2\ (m/s^2)"


"a_2=F_2/m_2=800/320=2.5\ (m/s^2)"


b)


"s_1=a_1t^2/2=2\cdot3^2/2=9\ (m)"


"s_2=a_2t^2/2=2.5\cdot3^2/2=11.25\ (m)" .Answer